Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Combo

Expert replies
by yellowho » Thu Jan 20, 2011 10:50 pm
If a fair die with six sides is tossed 6 times, what is the probability that at least two of outcomes will
be the same?
Join the discussion
Source: — Problem Solving |

by Everest » Thu Jan 20, 2011 11:05 pm
If a fair die with six sides is tossed 6 times, what is the probability that at least two of outcomes will
be the same?

probability (atleast 2) = 1 - 1/6 = 5/6
Join the discussion

by Rahul@gurome » Fri Jan 21, 2011 2:29 am
yellowho wrote:If a fair die with six sides is tossed 6 times, what is the probability that at least two of outcomes will be the same?
For this problem, it is easier to calculate the complement probability of the event rather than the probability of the event itself.

Probability that at least two outcomes are same = 1 - (Probability that all the outcomes are different)

Now, if we roll a six-sided dice 6 times, all the outcomes will be different if all possible outcomes occur, i.e. 1, 2, 3, 4, 5, and 6 in different rolls. Thus number of different sets of all different outcomes is nothing but the number of possible different arrangements of this 6 outcomes.

Thus, number of possible ways such that all the outcomes are different = 6! and total number of possible outcomes = 6^6

Probability that all the outcomes are different = 6!/(6^6)

Required probability = 1 - [6!/(6^6)]
Rahul Lakhani
Quant Expert
Gurome, Inc.
https://www.GuroMe.com
On MBA sabbatical (at ISB) for 2011-12 - will stay active as time permits
1-800-566-4043 (USA)
+91-99201 32411 (India)
Join the discussion

by Night reader » Fri Jan 21, 2011 3:25 am
Rahul@gurome wrote:
yellowho wrote:If a fair die with six sides is tossed 6 times, what is the probability that at least two of outcomes will be the same?
For this problem, it is easier to calculate the complement probability of the event rather than the probability of the event itself.

Probability that at least two outcomes are same = 1 - (Probability that all the outcomes are different)

Now, if we roll a six-sided dice 6 times, all the outcomes will be different if all possible outcomes occur, i.e. 1, 2, 3, 4, 5, and 6 in different rolls. Thus number of different sets of all different outcomes is nothing but the number of possible different arrangements of this 6 outcomes.

Thus, number of possible ways such that all the outcomes are different = 6! and total number of possible outcomes = 6^6

Probability that all the outcomes are different = 6!/(6^6)

Required probability = 1 - [6!/(6^6)]
quite scientific, whereas the smallest probability here is 1/6. Two events will be the same if the probability is not 1/6 -> this is included with our condition of 'at least 2 the same'. Hence P(2 the same)=1-Probability(not all the same) or 5/6
Join the discussion

by yellowho » Fri Jan 21, 2011 8:57 pm
Is there a different to do this? The OA is 319/324. It's exactly the same answer but that expression suggests that there's a different methodology. 6!/6^6 is quite the calculation.

[quote="Rahul@gurome"][quote="yellowho"]If a fair die with six sides is tossed 6 times, what is the probability that at least two of outcomes will be the same?[/quote]

For this problem, it is easier to calculate the complement probability of the event rather than the probability of the event itself.

Probability that at least two outcomes are same = 1 - (Probability that all the outcomes are different)

Now, if we roll a six-sided dice 6 times, all the outcomes will be different if all possible outcomes occur, i.e. 1, 2, 3, 4, 5, and 6 in different rolls. Thus number of different sets of all different outcomes is nothing but the number of possible different arrangements of this 6 outcomes.

Thus, number of possible ways such that all the outcomes are different = 6! and total number of possible outcomes = 6^6

Probability that all the outcomes are different = 6!/(6^6)

Required probability = 1 - [6!/(6^6)][/quote]
Join the discussion

by Rahul@gurome » Fri Jan 21, 2011 11:37 pm
yellowho wrote:Is there a different to do this? The OA is 319/324. It's exactly the same answer but that expression suggests that there's a different methodology. 6!/6^6 is quite the calculation.
I believe the method is same as I explained.
The calculation is not at all complicated as it looks. Proceed as follows,
  • 6!/6^6 = (1*2*3*4*5*6)/(6^6) = (4*5*6*6)/(6^6) = (4*5)/(6^4) = 5/(3*3*6*6) = 5/324
Hence, required probability = 1 - (6!/6^6) = 1 - 5/324 = 319/324
Rahul Lakhani
Quant Expert
Gurome, Inc.
https://www.GuroMe.com
On MBA sabbatical (at ISB) for 2011-12 - will stay active as time permits
1-800-566-4043 (USA)
+91-99201 32411 (India)
Join the discussion